2 a place progressives' has the following equation y=5 sin (100fy t_0.4x) calculate (1)amplitude (2)angular velocity (3)wave length (4)frequency and period (5)linear velocity
1
Expert's answer
2020-11-23T10:16:00-0500

solution

given equation

equation y=5 sin (100fy t_0.4x) comparing with standard equation

y=y0sin ("\omega" t-k x)

so

1)amplitude is y0=5m

2)angular velocity "\omega=100\times 3.14=314rad/sec"

3)wavelength

"\lambda=\frac{2\pi}{k}=\frac{2\times3.14}{0.4}=15.7m"

4)frequency

"\omega=2\pi f"

"f=\frac{\omega}{2\pi}=\frac{100\pi}{2\pi}=50Hz"


period "T=\frac{1}{f}==\frac{1}{50}=0.02 sec"

5)linear velocity

"v=r\omega=y_0\omega=5\times100\times3.14=1570m/s"